A box contains 12 red balls and 8 white balls. Two balls are drawn in succession. Find the following probabilities if the balls are drawn with replacement.
a.) A red ball is drawn then another red ball is drawn
b.) A red ball then a white ball
c.) Two white ball is drawn
d.) Both balls are of the same color

Answer by flame8855(424)

total placement = 20*20=400
1)12*12= 144 P= 144/200
2) 12*8=96, P=96/400
3)8*8=64 P = 64/400
4) so either white or red
P= (144+64)/400=208/400